Tohban Report 2011-07-20

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 13 July 2011
End Date: 20 July 2011
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Hazel Bain
List all reports


Solar Activity

It was another very quiet week. There are nine active regions on the disk, but few flares. The GOES event list shows 22 B and 2 C flares. RHESSI caught 16 (B) and 2 (C) of these, with 13 (B) and 1 (C) with excellent coverage. This level of activity is likely to continue over the next few days/week.

Memory Management

During the eclipses, data was taken for 4 minutes at start and end (compared to the entire eclipse the week before). The SSR reached a max fill of 46% and is trending downward since the eclipse data collection change.

Data Gaps

There were no data gaps this week.

Spacecraft Status

Detector 6 showed a couple arcs in the previous week's data. Anomalies were seen in flare spectra, where Detector 6 showed spikes up to 50 keV whereas the others did not. In response to this issue, a voltage stepdown of 20 steps (400V) took place on Tuesday (July 19, 20:12:37 UT) for detector 6, which remains segmented. At 2465 V, this is now the lowest-voltage detector. Detector 8 displayed some noise around 0430 on July 18, but the problem was not as bad as that noted in the last tohban report.

Personal tools